# Optimizing Multi-Stuttered Speech Classification: Leveraging Whisper’s Encoder for Efficient Parameter Reduction in Automated Assessment

### v2

@@ Abstract
The automated classification of stuttered speech has significant implications for timely assessments providing assistance to speech language pathologists. Despite notable advancements in the field, the cases in which multiple disfluencies occur in speech require attention. We have taken a progressive approach to fill this gap by classifying multi-stuttered speech more efficiently. The problem has been addressed by firstly curating a dataset of multi-stuttered disfluencies from SEP-28k audio clips. Secondly, employing Whisper, a state-of-the-art speech recognition model has been leveraged by using its encoder and taking the problem as multi label classification. Thirdly, using a 6 encoder layer Whisper and experimenting with various layer freezing strategies, a computationally efficient configuration of the model was identified. The proposed configuration achieved micro, macro, and weighted F1-scores of 0.88, 0.85, and 0.87, correspondingly on an external test dataset i.e. Fluency-Bank. In addition, through layer freezing strategies, we were able to achieve the aforementioned results by fine-tuning a single encoder layer, consequently, reducing the model’s trainable parameters from 20.27 million to 3.29 million. This research study unveils the contribution of the last encoder layer in the identification of disfluencies in stuttered speech. Consequently, it has led to a computationally efficient approach which makes the model more adaptable for various dialects and languages.
@@
Multi-stutter , Speech Classification , Whisper , Transformer , Deep Learning , Wav2vec2.0 , SEP-28k , Fluency Bank
PACS:
$0$ , $1$
MSC:
$0$ , $1$
@@
[inst1]organization=School of Electrical Engineering and Computer Science (SEECS),addressline=National University of Sciences and Technology (NUST), city=Islamabad, postcode=44000, country=Pakistan
[…]
[inst2]organization=College of Computer and Information Sciences (CCIS),addressline=Prince Sultan University, city=Riyadh, postcode=12435, country=Saudi Arabia
@@ 1 Introduction
Stuttering, a fluency disorder is a disruption in the normal flow of speech [1]. People with this disorder exhibit various types of disfluencies in their speech  [^1] https://www.asha.org/public/speech/disorders/stuttering/. The most common disfluencies include blocks, interjections, sound repetition, word repetition, and prolongation [1] which are demonstrated in Figure 1. Stuttering affects approximately 1% of the global population [2].
In this era of Artificial Intelligence, automated systems are assisting professionals in various domains i.e. medical, finance, education etc. Nevertheless, in the field of psychology, particularly for Speech Language Pathologists (SLP), the assessment of stuttering patients still requires considerable cognitive effort and is time-intensive. Conventionally, SLPs have to manually identify the disfluencies in speech, count them and calculate parameters like speech rate, stuttered words per minute etc [1]. Consequently, the outcome is dependent on the expertise of SLPs. This creates a need for an automated tool that can assist SLPs in evaluating stuttering patients.
@@
Although there have been numerous studies in this area, it still remains an evolving field. Since audio data is the most accessible source, most of the studies leverage it for research in the automated classification of disfluencies in stuttered speech [3, 4]. Lately, deep learning, specifically transformer-based models has emerged as a suitable choice for this task due to the complex nature of stuttered speech [5, 6]. Whisper [7], a weakly supervised transformer-based model developed by Open AI  [^2] Open AI Whisper: https://openai.com/research/whisper has proven to be a more resource-efficient and generalized model for automated classification of stuttered speech [8]. However, the study presented by [8] only deals with single stuttering, whilst in most real world scenarios, more than one disfluencies can co-occur. Although, studies presented by [9, 10] have addressed this issue by acquiring Wav2vec2.0 [11] and framing it as a multi-label classification for handling multi-stuttered speech instances. These studies lack in terms of speaker-exclusive dataset splits, high-annotator agreements, enough multi-stuttered data, and data quality.
This study aims to address the aforementioned gaps, and additionally, it also focuses on reducing the trainable parameters and finding the optimized configurations of the model. The goal is to have a resource-efficient solution and a generalized model for the multi-stuttered use cases. Following are the contributions of our study:
  • We have curated a multi-stuttered dataset using an open source dataset known as Stuttering Events in Podcasts (SEP-28k) and FluencyBank. The speakers were labeled manually, and then the audio of the same speakers was strategically concatenated to make it multi-stuttered.
  • The primary objective of this study is the automated classification of multi-stuttered disfluencies. This aspect contributes to the advancement of the field in real-world scenarios, and for that, we have leveraged the encoder-only Whisper model.
  • By limiting the learnable parameters of the model without compromising the performance of the model, we were able to optimize model efficiency, making it more suitable for real-world application. Specifically, we froze parameters including the feature extractor and encoder layers of Whisper.
@@ 2 Related Work
Disfluencies in stuttered speech are intricate in nature and have been addressed through various techniques applied to audio data. The acquisition of transformer-based models has revolutionized multiple domains, including computer vision [12], natural language processing [13], and speech processing [14]. In terms of automated stuttered speech, researchers have explored the transformer-based model Wav2vec2.0 which has shown commendable performances in various studies [15],[16],[17]. The labeled datasets SEP-28k and FluencyBank proposed by [18], has been acquired by the researchers. However, the majority of the studies are on single disfluency types, despite the fact that multiple disfluencies can occur in a single audio clip as well. In this section, we will cover the studies addressing multi-stutter labels.
@@
Disfluencies often co-occur with other types of disfluencies in real-world scenarios. Mapping this issue into deep learning parlance, each instance may consist of multiple labels, framing this as a multi-label classification problem [19]. In the domain of speech processing, numerous researchers have employed multi-label classification for emotion recognition [20], voice disorders [21] and speech recognition use-cases [22]. In the context of multi-stuttered scenarios, the authors [9] have conceptualized the issue as a multi-label classification problem. They have proposed Resnet50 as a solution for converting audio into spectrograms and feeding the images in the network. The researchers have utilized SEP-28k and FluencyBank datasets for training and testing. While the paper acknowledges the issue regarding data imbalance, it doesn’t explicitly address the resolution of the problem. In addition, the splits of the datasets play a major role in the outcome of the model, however, they are not discussed in the paper. The influence of dataset splits has been explored in the study [23], as it highlights the capability of the model’s efficacy.
@@
The authors [10] have also taken it as a multi-label problem and explored the cross-language capabilities of a model so it can be more generalized. In contrast to the aforementioned study [9], this study explicitly mentions the training, validation, and testing splits of the datasets. This study uses the SEP-28k Extended dataset, and KsoF dataset which is based on English and German stuttered speech. The proposed method uses Wav2vec2.0 along with a sequence classifier for end-to-end stuttering detection and classification. Furthermore, they have applied multi task learning to incorporate regularization. The evaluation metric acquired was F1-score for different combinations of training data and modifications. For the SEP-28k-E test set the F1-scores Block, Interjection, Prolongation, Sound Repetition, and Word Repetition achieved were 0.32, 0.77, 0.54, 0.50, and 0.56, respectively. Furthermore, for the fluency bank dataset, the F1-scores were 0.31, 0.82, 0.57, 0.61, 0.45 in correspondence with the above mentioned disfluency types. This has provided a new direction for solving the task at hand, however, the paper does not cater to the quality of the data by omitting the discussions on inter-annotator agreements in the labeling of these disfluencies. In Contrast, the study proposed by [15] has produced better results by only acquiring the high inter-annotator agreement labels. They have only considered the instances where all three annotators agreed on the labels. So, even if the quantity of the data was decreased, better results were achieved due to the quality of the dataset. In addition, they have not applied any preprocessing steps to the audio clips.
@@
The significance of preprocessing has been established in the paper [8]. The study has followed a series of data cleaning and preprocessing steps with significant improvement in the results. Furthermore, the approach proposed by [15] regarding the data quality of quantity was also used in this paper. The dataset used in this study is SEP-28k for training and FluencyBank for testing purposes. The paper acquires Whisper for the classification of disfluencies in stuttered speech, demonstrating superior performance compared to results obtained with the Wav2vec2.0 model. The paper has experimented with various data-splits achieving the highest average F1-score of 0.81. However, the study does not consider the multi stuttering cases.
@@
Hence, there exists a notable gap in the literature which necessitates the integration of multi-label classification and considers the proper dataset splits of training, validation and testing sets. Next, with the inclusion of preprocessing and cleaning of the input data based on transformer-based models, this research can be taken forward for real-world scenarios. However, another important aspect that needs to be addressed is the multi-label data for stuttered speech. While SEP-28k and FluencyBank encompass certain occurrences of multi-label instances. On the other hand, by prioritizing the dataset quality and choosing to include only instances where all three annotators agree on the same labels, the multi-label number of instances was reduced significantly. In this context, scarcity presents a challenge to the learning of the models. Therefore, this emphasizes the need for a dataset that specifically is tailored to multi-stuttered speech. Lastly, building upon the findings presented in the referenced paper [8], where Whisper has outperformed Wav2vec2.0. Our study, in conjunction with other mentioned gaps, will also leverage Whisper for the automated classification of disfluencies in the context of multi-stuttered speech.
@@ 3.1 Datasets
In this study, two primary datasets SEP-28k [18] and FluencyBank [24] have been leveraged for training and testing the models. SEP-28k consists of  28k instances, and each audio clip is 3 seconds long. These clips were extracted from eight podcasts and labeled with five major disfluency types; prolongation, sound repetition, word repetition, interjections, and blocks. Similarly, FluencyBank has been labelled and consists of the same audio length as SEP-28k. It consists of  4k instances and is mostly leveraged for testing purposes. Lastly, both datasets consist of audio of English speakers.
To improve the dataset’s quality, data cleaning steps from the paper [8] have been followed considering their improved outcomes. At first, the experimentation was limited to cases where all three annotators agreed on a certain label. Next, labels that accounted for less than 1% of the dataset’s representation were eliminated. These were non-disfluent labels, like Natural Pause, Hard to Understand, Speechless, Bad Audio Quality, and Music. Moreover, in order to maintain uniformity, audio fragments that lasted less than three seconds were also not included. To reduce the possibility of influencing the outcome of the model, background noises were reduced from the audio clips. Finally, audio samples that included multiple speakers were also excluded as they may complicate classification and interpretation.
In the study presented by [23], the influence of data partitioning has been explored. The authors have proposed and experimented with three speaker-exclusive dataset splits i.e. SEP-28k-E, SEP-28k-E-T and SEP-28k-D. The training split of SEP-28k-E consists of clips with the top four dominant speakers. The remaining clips are partitioned for testing and validation. SEP-28k-T and SEP-28k-D are similar as they are using the testing and validation split of SEP-28k-E in their training splits, and testing them on the top four dominant speakers. The splits of SEP-28k-E can exhibit the capability of a model by training them on a few speakers with more audio clips and testing on various speakers. On the other hand, SEP-28k-T and SEP-28k-D can identify the ability of the model to learn from fewer examples of various speakers. In the experimental work of [8], the authors have investigated a different setup that combines training and validation data, using the test split for validation purposes. These datasets, which we will refer to as SEP-28k-E-merged and SEP-28k-T-merged, are shown in Table 1 with their splits. These additional splits are also included in the respective study.
Table 1: Data Split Configurations for Experimental Study
@@ 3.2 Data Curation for Multi-Stutter
Therefore, to curate multi-stuttered data, specific conditions needed to be met. The audio from the same episodes and the same speakers were required. Since in each episode featured multiple speakers, we manually had to label each of them so they could be concatenated. Next, two audio samples consisting of distinct disfluent labels were then concatenated using Python library AudioSegment. In addition, if both the labels were ”NoStutteredWords”, they were also concatenated. The pipeline for the multi-stuttered dataset is illustrated in Figure 2.
Figure 2: Pipeline for Curation of Multi-stuttered data: Condition 1 involves acquiring audio from the same episodes, Condition 2 requires audios featuring the same speakers, and Condition 3, proceeds with concatenation if two audios have distinct disfluency labels or both have the label ”NoStutteredWords”.

### v3

# Conceptualization of this study, Methodology, Software

@@ Abstract
The automated classification of stuttered speech has significant implications for timely assessments providing assistance to speech language pathologists. Despite notable advancements in the field, the cases in which multiple disfluencies occur in speech require attention. We have taken a progressive approach to fill this gap by classifying multi-stuttered speech more efficiently. The problem has been addressed by firstly curating a dataset of multi-stuttered disfluencies from open source dataset SEP-28k audio clips. Secondly, employing Whisper, a state-of-the-art speech recognition model has been leveraged by using its encoder and taking the problem as multi label classification. Thirdly, using a 6 encoder layer Whisper and experimenting with various layer freezing strategies, a computationally efficient configuration of the model was identified. The proposed configuration achieved micro, macro, and weighted F1-scores of 0.88, 0.85, and 0.87, correspondingly on an external test dataset i.e. Fluency-Bank. In addition, through layer freezing strategies, we were able to achieve the aforementioned results by fine-tuning a single encoder layer, consequently, reducing the model’s trainable parameters from 20.27 million to 3.29 million. This research study unveils the contribution of the last encoder layer in the identification of disfluencies in stuttered speech. Consequently, it has led to a computationally efficient approach, 83.7% less parameters to train, making the proposed approach more adaptable for various dialects and languages.
@@
{highlights}
Curated a multi-stuttered disfluencies using SEP-28k and FluencyBank datasets, enhancing them by manually labeling speakers and strategically concatenating audio to simulate multi-stuttered speech.
Employed the Whisper model’s encoder and reframed the problem as a multi-label classification for multi-stuttered speech. Achieved high micro, macro, and weighted F1-scores of 0.88, 0.85, and 0.87, respectively on an external test dataset, demonstrating significant advancements in handling multiple disfluencies.
Achieved a substantial reduction in trainable parameters from 20.27 million to 3.29 million by employing strategic layer freezing, resulting in a computationally efficient model adaptable to various dialects and languages.
@@ 1 Introduction
Stuttering, a fluency disorder is a disruption in the normal flow of speech Shipley and McAfee, (2019). People with this disorder exhibit various types of disfluencies in their speech  [^1] https://www.asha.org/public/speech/disorders/stuttering/. The most common disfluencies include blocks, interjections, sound repetition, word repetition, and prolongation Shipley and McAfee, (2019) which are demonstrated in Figure 1. Stuttering affects approximately 1% of the global population Yairi and Ambrose, (2013).
In this era of Artificial Intelligence, automated systems are assisting professionals in various domains i.e. medical, finance, education etc. Nevertheless, in the field of psychology, particularly for Speech Language Pathologists (SLP), the assessment of stuttering patients still requires considerable cognitive effort and is time-intensive. Conventionally, SLPs have to manually identify the disfluencies in speech, count them and calculate parameters like speech rate, stuttered words per minute etc Shipley and McAfee, (2019). Consequently, the outcome is dependent on the expertise of SLPs. This creates a need for an automated tool that can assist SLPs in evaluating stuttering patients.
@@
Although there have been numerous studies in this area, it still remains an evolving field. Since audio data is the most accessible source, most of the studies leverage it for research in the automated classification of disfluencies in stuttered speech Almasoud et al., (2022); Sheikh et al., (2021). Lately, deep learning, specifically transformer-based models has emerged as a suitable choice for this task due to the complex nature of stuttered speech Bayerl et al., 2023a ; Sebastian et al., (2022). Whisper Radford et al., (2023), a weakly supervised transformer-based model developed by Open AI  [^2] Open AI Whisper: https://openai.com/research/whisper has proven to be a more resource-efficient and generalized model for automated classification of stuttered speech Ameer et al., (2023). However, the study presented by Ameer et al., (2023) only deals with single stuttering, whilst in most real world scenarios, more than one disfluencies can co-occur. Although, studies presented by Jouaiti and Dautenhahn, (2022); Bayerl et al., 2023b have addressed this issue by acquiring Wav2vec2.0 Baevski et al., (2020) and framing it as a multi-label classification for handling multi-stuttered speech instances, still the domain needs more work.
1.1 Research Gaps and Objectives
Despite significant advancements in disfluency classification, several gaps remain in the classification of multi-stuttered disfluencies. The aforementioned studies lack in terms of speaker-exclusive dataset splits, high annotator agreements, enough multi-stuttered data, and data quality. This study aims to address these gaps, and additionally, it also focuses on reducing the trainable parameters and finding the optimized configurations of the model. The goal is to have a resource-efficient solution and a generalized model for multi-stuttered use cases.
1.1.1 Curating a Comprehensive Multi-Stuttered Dataset:
Objective: To curate a quality dataset of multi-stuttered speech by combining and enhancing existing open-source datasets.
Contribution: We have curated a multi-stuttered dataset using an open-source dataset known as Stuttering Events in Podcasts (SEP-28k) and FluencyBank. The speakers were labeled manually, and then the audio of the same speakers was strategically concatenated to create multi-stuttered samples.
1.1.2 Automated Classification of Multi-Stuttered Disfluencies:
Objective: To implement a strategy for the classification of multi-stuttered disfluencies.
Contribution: The primary objective of this study is the automated classification of multi-stuttered disfluencies. This aspect contributes to the advancement of the field in real-world scenarios, and for that, we have leveraged the encoder-only Whisper model.
1.1.3 Model Optimization for Real-World Applications:
Objective: To optimize the model without sacrificing classification performance.
Contribution: By limiting the learnable parameters of the model without compromising its performance, we optimized model efficiency, making it more suitable for real-world applications. Specifically, we froze parameters including the feature extractor and encoder layers of Whisper.
@@ 2 Related Work
Disfluencies in stuttered speech are intricate in nature and have been addressed through various techniques applied to audio data. The acquisition of transformer-based models has revolutionized multiple domains, including computer vision Parvaiz et al., (2023), natural language processing Patwardhan et al., (2023), and speech processing Latif et al., (2023); Nowakowski et al., (2023). In terms of automated stuttered speech, researchers have explored the transformer-based model Wav2vec2.0 which has shown commendable performances in various studies Mohapatra et al., (2022), Bayerl et al., 2022b , Bayerl et al., 2022a , Al-Banna et al., (2024). The labeled datasets SEP-28k and FluencyBank proposed by Lea et al., (2021), has been acquired by the researchers. However, the majority of the studies are on single disfluency types, despite the fact that multiple disfluencies can occur in a single audio clip as well. In this section, we will cover the studies addressing multi-stutter labels.
@@
Disfluencies often co-occur with other types of disfluencies in real-world scenarios. Mapping this issue into deep learning parlance, each instance may consist of multiple labels Tsoumakas and Katakis, (2007) , framing this as a multi-label classification problem. In the domain of speech processing, numerous researchers have employed multi-label classification for emotion recognition Slimi et al., (2022), voice disorders Junior et al., (2023) and speech recognition use-cases Hossain et al., (2022). In the context of multi-stuttered scenarios, the authors Jouaiti and Dautenhahn, (2022) have conceptualized the issue as a multi-label classification problem. They have proposed ResNet-50 as a solution for converting audio into spectrograms and feeding the images in the network. The researchers have utilized SEP-28k and FluencyBank datasets for training and testing. While the paper acknowledges the issue regarding data imbalance, it doesn’t explicitly address the resolution of the problem. In addition, the splits of the datasets play a major role in the outcome of the model, however, they are not discussed in the paper. The influence of dataset splits has been explored in the study Bayerl et al., 2022c , as it highlights the capability of the model’s efficacy.
@@
The authors Bayerl et al., 2023b and Wagner et al., (2024) have also taken it as a multi-label problem and explored the cross-language capabilities of a model so it can be more generalized. In contrast to the aforementioned study Jouaiti and Dautenhahn, (2022), this study explicitly mentions the training, validation, and testing splits of the datasets. This study uses the SEP-28k Extended dataset, and KSoF dataset which is based on English and German stuttered speech. The proposed method uses Wav2Vec2.0 along with a sequence classifier for end-to-end stuttering detection and classification. Furthermore, they have applied multi task learning to incorporate regularization. The evaluation metric acquired was F1-score for different combinations of training data and modifications. For the SEP-28k-E test set the F1-scores of Block, Interjection, Prolongation, Sound Repetition, and Word Repetition achieved were 0.32, 0.77, 0.54, 0.50, and 0.56, respectively. Furthermore, for the FluencyBank dataset, the F1-scores were 0.31, 0.82, 0.57, 0.61 and 0.45 in correspondence with the above mentioned disfluency types. This has provided a new direction for solving the task at hand, however, the paper does not cater to the quality of the data by omitting the discussions on inter-annotator agreements in the labeling of these disfluencies. In Contrast, the study proposed by Mohapatra et al., (2022) has produced better results by only acquiring the high inter-annotator agreement labels. They have only considered the instances where all three annotators agreed on the labels. So, even if the quantity of the data was decreased, better results were achieved due to the quality of the dataset. In addition, they have neither applied any preprocessing steps to the audio clips nor catered multi-stuttered use-case scenario.
@@
The significance of preprocessing has been established in the paper Ameer et al., (2023). The study has followed a series of data cleaning and preprocessing steps with significant improvement in the results. Furthermore, the approach proposed by Mohapatra et al., (2022) regarding the data quality of quantity was also used in this paper. The dataset used in this study is SEP-28k for training and FluencyBank for testing purposes. The paper acquires Whisper for the classification of disfluencies in stuttered speech, demonstrating superior performance compared to results obtained with the Wav2Vec2.0 model. The paper has experimented with various data-splits achieving the highest average F1-score of 0.81. However, the study does not consider the multi stuttering cases.
@@
Hence, there exists a notable gap in the literature which necessitates the integration of multi-label classification and considers the proper dataset splits of training, validation and testing sets. Next, with the inclusion of preprocessing and cleaning of the input data based on transformer-based models, this research can be taken forward for real-world scenarios. However, another important aspect that needs to be addressed is the multi-label data for stuttered speech. While SEP-28k and FluencyBank encompass certain occurrences of multi-label instances. On the other hand, by prioritizing the dataset quality and choosing to include only instances where all three annotators agree on the same labels, the multi-label number of instances was reduced significantly. In this context, scarcity presents a challenge to the learning of the models. Therefore, this emphasizes the need for a dataset that specifically is tailored to multi-stuttered speech. Lastly, building upon the findings presented in the referenced paper Ameer et al., (2023), where Whisper has outperformed Wav2Vec2.0. Our study, in conjunction with other mentioned gaps, will also leverage Whisper for the automated classification of disfluencies in the context of multi-stuttered speech.
@@ 3.1 Datasets
In this study, two primary datasets SEP-28k Lea et al., (2021) and FluencyBank Ratner and MacWhinney, (2018) have been leveraged for training and testing the models. SEP-28k consists of  28k instances, and each audio clip is 3 seconds long. These clips were extracted from eight podcasts and labeled by experts with five major disfluency types; prolongation, sound repetition, word repetition, interjections, and blocks Lea et al., (2021). Similarly, in the paper Lea et al., (2021), FluencyBank has also been labelled and consists of the same audio length as SEP-28k. It consists of  4k instances and is mostly leveraged for testing purposes. Lastly, both datasets consist of audio of English speakers. In addition, in this study, SEP-28k have been used for training, validation and internal testing. FluencyBank has been used for the external testing of the model. In the study presented by Bayerl et al., 2022c , the influence of data partitioning has been explored. The authors have proposed and experimented with three speaker-exclusive dataset splits i.e. SEP-28k-E, SEP-28k-E-T and SEP-28k-D. The training split of SEP-28k-E consists of clips with the top four dominant speakers. The remaining clips are partitioned for testing and validation. SEP-28k-T and SEP-28k-D are similar as they are using the testing and validation split of SEP-28k-E in their training splits, and testing them on the top four dominant speakers. The splits of SEP-28k-E can exhibit the capability of a model by training them on a few speakers with more audio clips and testing on various speakers. On the other hand, SEP-28k-T and SEP-28k-D can identify the ability of the model to learn from fewer examples of various speakers. In the experimental work of Ameer et al., (2023), the authors have investigated a different setup that combines training and validation data, using the test split for validation purposes. These datasets, which we will refer to as SEP-28k-E-merged and SEP-28k-T-merged, are shown in Table 1 with their splits. These additional splits are also included in the respective study.
To improve the dataset’s quality, data cleaning steps from the paper Ameer et al., (2023) have been followed considering their improved outcomes. At first, the experimentation was limited to cases where all three annotators agreed on a certain label. Next, labels that accounted for less than 1% of the dataset’s representation were eliminated. These were non-disfluent labels, like Natural Pause, Hard to Understand, Speechless, Bad Audio Quality, and Music. Moreover, in order to maintain uniformity, audio fragments that lasted less than three seconds were also not included. To reduce the possibility of influencing the outcome of the model, background noises were reduced from the audio clips. Finally, audio samples that included multiple speakers were also excluded as they may complicate classification and interpretation.
Table 1: Split Configurations for Experimental Study
@@ 3.2 Data Curation for Multi-Stutter
Therefore, to curate multi-stuttered data, specific conditions needed to be met. The audio from the same episodes and the same speakers were required. Since in each episode featured multiple speakers, we manually had to label each of them so they could be concatenated. Next, two audio samples consisting of distinct disfluent labels were then concatenated using Python library AudioSegment. In addition, if both the labels were "NoStutteredWords", they were also concatenated. The pipeline for the multi-stuttered dataset is illustrated in Figure 2.
Figure 2: Pipeline for Curation of Multi-stuttered data: Condition 1 involves acquiring audio from the same episodes, Condition 2 requires audios featuring the same speakers, and Condition 3, proceeds with concatenation if two audios have distinct disfluency labels or both have the label "NoStutteredWords".